OpenStudy (anonymous):

Determine the number and type of solutions for 5x^2 + 1x + 7 = 0. 1 Rational Double 2 Irrational 2 Rational 2 Complex

OpenStudy (anonymous):

these are the answers given to do the mc

OpenStudy (lalaly):

2 complex

OpenStudy (anonymous):

2 complex

OpenStudy (anonymous):

Possible intermediate steps: 5 x^2+x+7 = 0 Solve the quadratic equation by completing the square: Divide both sides by 5: x^2+x/5+7/5 = 0 Subtract 7/5 from both sides: x^2+x/5 = -7/5 Add 1/100 to both sides: x^2+x/5+1/100 = -139/100 Factor the left hand side: (x+1/10)^2 = -139/100 Take the square root of both sides: abs(x+1/10) = (i sqrt(139))/10 Eliminate the absolute value: x+1/10 = -(i sqrt(139))/10 or x+1/10 = (i sqrt(139))/10 Subtract 1/10 from both sides: x = -1/10 i (-i+sqrt(139)) or x+1/10 = (i sqrt(139))/10 Subtract 1/10 from both sides: x = -1/10 i (-i+sqrt(139)) or x = 1/10 i (i+sqrt(139)) so both are of type , complex